Strawberrys:

-\frac{b}{5}-27= − 5 b ​ −27= \,\,-21 −21

Frosted:

Can you try reposting the question and/or post a screenshot of your question?

Strawberrys:

yes

Strawberrys:

Ferrari:

Ok

Ferrari:

First, add 39 to both sides of the equation

Strawberrys:

i did that

Ferrari:

3b − 39 + 39 = −78 + 39

Ferrari:

3b = −39

Frosted:

Okay, I guess you can just take my question... e.e

Ferrari:

Divide both sides of the equation by 3.

Ferrari:

\[\frac{ 3b }{ 3 } = \frac{ -39 }{ 3 }\]

Ferrari:

Solve this problem, and then you will discover that: b = ...

Ferrari:

Dividing the sides by 3 is the final step.

Ferrari:

Okay, I guess I will need to do it. b = -13.

Strawberrys:

thx i got it